OpenStudy (anonymous):

The volume of a cone is three hundred cubic kilometers. If the radius of the base is 5 kilometers, what is the height of the cone? Use three point one four for pi. Show your work and round to the nearest tenth.

OpenStudy (anonymous):

300(3)/pi(r^2)

OpenStudy (anonymous):

286.624203822/5^2

OpenStudy (anonymous):

Height=11.5

OpenStudy (whpalmer4):

volume of a cone: \[V=\frac{1}{3}\pi r^2h\]Solve for h\[\frac{3V}{\pi r^2} = h\]Plug in the numbers:\[\frac{3*300}{3.14*5*5} = 11.4592\]round to the nearest 10th gives you 11.5